Suppose that the pre-tax price of cigarettes is $3.00 per pack and the post-tax price of cigarettes

is $4.00 per pack. The tax is $1.50 per pack and is paid by the retailers of cigarettes to the

government. Six hundred packs were sold prior to the tax increase and 500 packs were sold

after the tax. What is the consumer tax burden? (5 + 5 = 10)


1
Expert's answer
2021-02-04T14:42:55-0500
"600\times(4-3)+500\times 1.5=1350"